OpenStudy (ahsome):

Find the horizontal force required to keep a small box of mass 200g at rest on a smooth plane inclined at 30 degrees to the horizontal gardiner

OpenStudy (ahsome):

@shamim?

OpenStudy (ahsome):

Forget the gardiner

OpenStudy (shamim):

Csn u draw it???

OpenStudy (ahsome):

 I think

OpenStudy (shamim):

U know the weight w=mg=0.2*9.8=?

OpenStudy (ahsome):

Yup. W=1.96N

OpenStudy (shamim):

Yes u r right!!!

OpenStudy (ahsome):

Now what? I found that Fp is 0.98N

OpenStudy (shamim):

Now component of w along inclined plane =1.96*sin30=?

OpenStudy (ahsome):

0.98N

OpenStudy (shamim):

So we need to apply 0.98N along the inclined plane up to make the object stationary

OpenStudy (shamim):

We hv to find out the horizontal force. Right?

OpenStudy (ahsome):

Yup

OpenStudy (shamim):

Let the horizontal applied force =F

OpenStudy (ahsome):

Ok

OpenStudy (shamim):

Component of F along the inclined plane is=0.98N

OpenStudy (ahsome):

Yes

OpenStudy (shamim):

So we can write F*cos30=0.98 F=?

OpenStudy (ahsome):

\[F=\frac{0.98}{cos(30)}\]?

OpenStudy (ahsome):

That works! THANK YOU. But why divide by cos(30)?

OpenStudy (shamim):

F*3=9 F=9/3 Right?

OpenStudy (shamim):

F*cos30=0.98 F=0.98/cos30 Right?

OpenStudy (ahsome):

Exactly. But why could we say F = 0.98*cos(30)? 

OpenStudy (ahsome):

Is it because the force itself is horizontal, its just that part acts up the plane?

OpenStudy (shamim):

F is along horizontal. So the inclined plane makes angle with horizontsl is=30 degree?

OpenStudy (ahsome):

Yes

OpenStudy (shamim):

Did u get it??!

OpenStudy (ahsome):

I see. Thanks alot :D

OpenStudy (shamim):

U r most welcome!!!!

OpenStudy (ahsome):

Fanned as well :). Got to go to sleep. Night :)

OpenStudy (shamim):

I m also enjoyed a lot!!!!!!!

OpenStudy (shamim):

Anywsy frm where u r???!! Usa??

OpenStudy (ahsome):

@shamim, Australia :D